Grade 12Mechanics

A stone is dropped from the top of a tower and one s later, a second stone is thrown vertically downward with a velocity 20m/s. The second stone will overtake the first stone after travelling a distance of
a.13m. b.15m. c.11.25m. d. 19.5m

Feyba Jijimon
9 Years agoGrade 12

1 Answer

Khimraj
9 Years ago
Write the displacement equation s1 = 1/2gt² and s2 = 20(t-1) +1/2g(t-1)² and as s1 = s2 so 20(t-1) + 1/2g(1-2t) =0 then 20t-20+5-10t =0 so t= 3/2 sec so distance is 1/2g(3/2)² =45/4.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer